OpenStudy (anonymous):

Amanda exercised for 10 minutes every day in the first week, 20 minutes in the second week, 30 minutes in the third week, and 40 minutes in the fourth week. Billy exercised for 5 minutes every day in the first week, 10 minutes in the second week, 20 minutes in the third week, and 40 minutes in the fourth week. Which statement best describes the methods used by Amanda and Billy to increase the time they spent exercising?

OpenStudy (anonymous):

(A) Amanda's method is linear because the number of minutes increased by an equal number every week. (B) Billy's method is linear because the number of minutes increased by an equal factor every week. (C) Both Billy's and Amanda's methods are exponential because the number of minutes increased by an equal factor every week. (D) Both Billy's and Amanda's methods are exponential because the number of minutes increased by an equal number every week.

OpenStudy (anonymous):

I need help on this question plz. @HelpBlahBlahBlah @hartnn @iambatman @kropot72 @KamiBug

OpenStudy (anonymous):

I know that Amanda has an arithmetic sequence with a difference of 10. Billy has a geometric sequence with a constant ratio of 2.

OpenStudy (anonymous):

But I don't know how those are described in the options above. -_-

OpenStudy (anonymous):

@Godlovesme can u help me plz?

OpenStudy (anonymous):

@Diyadiya @Directrix @Sadworld @Michele_Laino @One098 Can u guys plz help me out?

OpenStudy (anonymous):

Can you have 2 answers?

OpenStudy (anonymous):

no only 1; it's a multiple choice question on FLVS

OpenStudy (anonymous):

If not, the answer would be D since both increase exponentially since the number of minuets increase by an equal number each week. Amanda by 10 and Billy by 2(x)

OpenStudy (anonymous):

It can't be C see since Amanda's increase is not due to factors. A and B doesn't answer the question.

OpenStudy (anonymous):

Oh, it's C. I just did the math.

OpenStudy (michele_laino):

mins of Billy are: 5-10-20-40 can they mamatize with a linear or proportional function?

OpenStudy (michele_laino):

note that: 10=2*5 20=2*10 40=2*20 so? @happyrosy

OpenStudy (godlovesme):

linear function is if the pattern is the constant like 5,10,15,20 --> it increases with constant rate of 5 (+5) and exponential is multiplication like 5,10,20,40 --> x2 each value confused @happyrosy

OpenStudy (anonymous):

So the linear function is Amanda's sequence?

OpenStudy (godlovesme):

yes :) what about billy's?

OpenStudy (anonymous):

Billy's is exponential.

OpenStudy (godlovesme):

correct!!

OpenStudy (anonymous):

So B, C, and D are ruled out right? Making A the correct answer.

OpenStudy (michele_laino):

what is the xponential function of Billy?

OpenStudy (anonymous):

Billy's gets multiplied by 2 each time.

OpenStudy (michele_laino):

I think:\[5*2^{n}\] where n is the number of the week

OpenStudy (michele_laino):

n=0, 1,2,3

OpenStudy (anonymous):

Yeah that's another way because when n= 0, f(0) equals to 5

OpenStudy (anonymous):

when n= 1, f(1) equals to 10...n= 2, f(n)= 20..and so on

OpenStudy (michele_laino):

you are right! @happyrosy so I rewrite the function as below: \[5*2^{n-1}\] where n=1, 2, 3, 4

OpenStudy (anonymous):

okay so would A be the answer then?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

THANK YOU ALL FOR HELPING ME OUT!! :) I understood everything.

OpenStudy (michele_laino):

thank you! :) @happyrosy

OpenStudy (godlovesme):

anytime :) @happyrosy thanks for the medal

OpenStudy (anonymous):

np :)